OpenStudy (anonymous):

find the equation of the line that passes through the pair of points. write your answer in slope-intercept form. (2,7),(-2,-1)

OpenStudy (anonymous):

(y-y1)/(x-x1)=(y2-y1)/(x2-x1) (y-7)/(x-2)=-1-7/-2-2 (y-7)/(x-2)=-8/-4 y-7=2(x-2) y-7=2x-4 2x-y+3=0